Consider a fever thermometer that contains 0.01 cm³ of mercury, Find the diameter of the capillary if a 1 C° change corresponds to a level change of 0.5 cm . Assume the glass does noe expand.


1
Expert's answer
2022-02-02T17:34:43-0500

"h=h_0(1+\alpha T),"

"\Delta h=\alpha Th_0,"

"h_0=\frac{\Delta h}{\alpha T}=3.6~cm;"

"V=\frac{\pi d^2}4h_0,"

"d=\sqrt{\frac{4V}{\pi h_0}}=0.6~mm."